OpenStudy (anonymous):

How do you solve 2/3(x-5)^2=40? Thanks.

OpenStudy (anonymous):

Move 40 from the RHS to the LHS and then multiply the LHS by 3/2. \[\frac{3}{2} \left(\frac{2}{3} (x-5)^2-40\right)=0 \]Expand and simplify the above.\[x^2-10 x-35=0 \]\[x=5\pm 2 \sqrt{15} \]

OpenStudy (anonymous):

how did you get -35